What is the next two numbers in this sequence, 4,9,7,14,10,19,13,24,16,29
Answer by jim_thompson5910(35256)

Look at the odd numbered terms (in red)

4,9,7,14,10,19,13,24,16,29

Going from 4, to 7, to 10, etc tells us we're adding 3 to each term. This is because 4+3 = 7, 7+3 = 10, etc.

So part of the rule is this: if n is an odd number, then you use the rule: an = 3n+1 where n starts at n = 1. This rule only works if n is odd though.

If n is even, then we have this sequence (look at the above sequence, but instead of the red terms, focus on the black terms): 9, 14, 19, 24, 29

Now we're adding 5 to each term. So the rule is: if n is even, then use the formula an = 5n+4

So in the end, we have this rule

If n is odd, use the formula an = 3n+1, where n starts at n = 1
If n is even, use the formula an = 5n+4, where n starts at n = 2

That completes the entire sequence because any number is either even or odd.


The next term would be odd, so use the odd formula: 3n+1 = 3(6)+1 = 19. So the next term after 29 is 19

The term after that is even, so use 5n+4 = 5(6)+4 = 34.


So the next two terms are: 19 and 34

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------